Today is January 3. Your friend David has just bought a futures contract on a stock index, and the contract specifies one year to expiration. The current share price is $80, and the annually compounded interest rate is 10%. The stock will pay quarterly dividends of $2 during the next year, with dividends payments on the following dates:

  • January 25
  • April 25
  • July 25
  • October 25

Assume that this is a non-leap year.

a.       What is the futures price on this contract on January 3?

b.       What is the cost-of-carry on this futures contract on January 3?

c.       What is the value of this futures contract on February 17 if the spot price turns out to be $90 on that day?

d.       Suppose that, instead of paying a quarterly dividend of $2, the stock index has an annual dividend yield of 10%. The continuously compounded interest rate is 10.52%. What is the price of the index futures on January 3? What is its value on February 17 when the index turns out to be $90?
Answer #1

Answer a) The future price

F0 = (S0 – I)erT

I = PV(d1)+PV(d2)...+PV(d4) = d1e-rt1 +d2e-rt2 + d3e-rt3+d4e-rt4

Future price at 3 Jan,

Dividend Date Time(days) Time (year) e^rt PV(d)
2 25-Jan 22 days 0.060 1.01 2.01
2 25-Apr 90 days 0.247 1.02 2.05
2 25-Jul 91 days 0.249 1.03 2.05
2 25-Oct 91 days 0.249 1.03 2.05
Total (I) 8.16

F0 = (80-8.16)* e0.10*1 = $ 79.397 = $ 79.40.

Answer b) Futures price = Spot price + cost of carry

Cost of carry = Fo - So = 79.40-80 = (- 0.60)

Answer c) Value Future contract , f = (F0-K)*e-rt = (91.55-79.40)*e-0.1*0.877 = $ 11.13.

as K = $ 79.40 ,

at 17 Feb , F0 = (90 - 6.14)*e0.10*0.877 = 91.55

present value of Income at 17 feb , and time for contract = 365-45= 320 days (0.877 years)

Dividend Date Time(days) Time (year) e^rt PV(d)
0 25-Jan 22 days 0.060 1.01 0.00
2 25-Apr 67 days 0.184 1.02 2.04
2 25-Jul 91 days 0.249 1.03 2.05
2 25-Oct 91 days 0.249 1.03 2.05
Total (I) 6.14

Answer d) The price in future at 3 jan,  F0=S0e (r–y)T = 80*e(0.1052-0.10)*1= $ 80.42

At 17 Feb , f = (F0-K)*e-rt = (90.412-80.42)*e-0.1*0.877 = $ 9.16

K = $ 80.42

at 17 Feb , F0 = 90*e(0.1052-0.10)*0.877 =90.412